how much 25% of solution must be mixed with 70% solution to make 100ml of 58 solution
Answer by nerdybill(7384)

how much 25% of solution must be mixed with 70% solution to make 100ml of 58 solution
.
Let x = amount (ml) of 25% solution
then
100-x = amount of 70% solution
.
.25x + .70(100-x) = .58(100)
.25x + 70-.70x = 58
70-.45x = 58
-.45x = 58-70
-.45x = -12
x = 26.67 ml (amount of 25% solution)
.
amount of 70% solution:
100-x = 100-26.67 = 73.33 ml